OpenStudy (anonymous):

If the average of 10,20,30,40, and a is 50, what is the value of a??? plz help im in 6th grade :(((

OpenStudy (zarkon):

150

OpenStudy (anonymous):

(10+20+30+40+a)/5 = 50 (100+a)=50*5 100+a=250 a=250-100 a=150

OpenStudy (anonymous):

m not sure

OpenStudy (anonymous):

 a=150